A uniform solid sphere of radius R = 1.9 km produc
Answer #1

utside the sphere:

ag = GM/R12 with G = gravitation constant, M = mass of sphere, R1 = radius of sphere.

ag/4 = GM/R22 with R2 = unknown distance outside of the sphere.

R2/R1 = √{(4GM/(ag))/(GM/(ag)} = √4 =2

R2 = R1*2 = 1.9*2 km =3.8Km
-------------------------
2) inside the sphere:

ag decreases linearly with the radius, and is = 0 at the center.

So R = R1*1/4 = 1.9/4=0.475 km